Circular No. 7585

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001V IN NGC 3987 S. Jha, T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report the discovery of a supernova (mag 16) by P. Berlind on Feb. 19.38 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph). Concurrent images taken with the Observatory’s 1.2-m telescope (+ 4Shooter) yield the following precise position for SN 2001V: α = 11^{h}57^{m}24^{s}.93, δ = +25^{o}12′09′′.0 (equinox 2000.0, uncertainty ± 0′′.3), which is 52′′ east and 28′′ north of the host galaxy’s nucleus. The supernova spectrum exhibits a blue continuum with broad features, identifying this as a type-Ia supernova well before maximum light. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 4502 km/s for the host galaxy yields an expansion velocity for the supernova of 14 000 km/s, measured with the Si II (rest 635.5 nm) feature.

SUPERNOVA 2001T IN MCG –02-37-6 Matheson, Jha, Challis, and Kirshner also report that a spectrum of

SN 2001T (cf. IAUC 7581), obtained by Berlind (as above) on Feb. 19.49 UT, shows it to be a type-II supernova near maximum light. The spectrum consists of a blue continuum with superposed P-Cyg lines of hydrogen and Ca II. Adopting the NED recession velocity of 4167 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 7600 km/s.

COMET C/2001 C5 (SOHO) Further to IAUC 7582, D. Hammer reports his measurements for a

comet found by M. Oates on SOHO website images. C/2001 C5 was visible in both the C3 and C2 coronagraphs, and D. Biesecker provides the following post-perihelion V magnitudes from the C2 data (the C3 data being poor due to vignetting): Feb. 13.854, 5.4; 13.896, 5.0; 13.938, 5.0; 13.979, 4.9; 14.021, 4.9; 14.064, 4.9; 14.104, 5.3; 14.146, 5.6; 14.163, 6.3; 14.188, 7.4. The reduced observations and retrograde parabolic orbital elements (T = 2001 Feb. 13.3 TT, q = 0.026 AU, i = 166^{o}.3; a direct orbit solution is also possible) by B. G. Marsden appear on MPEC 2001-D07.

Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 C5 Feb. 13.721  21^{h}41.^{m}2  −13^{o}25′
2001 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green